# A compact Sr magneto-optical trap system for field-deployable optical lattice clocks

### v1

@@ Appendix A Dependence of the loss rate of trapped atoms on background gas pressure
where $R$ is the loading rate of trapped atoms and $\gamma$ denotes the loss coefficient due to collisions with background gas. The dominant component of the background gas is hydrogen ($\mathrm{H_{2}}$). In the following, we assume that the background gas consists solely of $\mathrm{H_{2}}$. Here, $\gamma$ is proportional to the partial pressure of hydrogen $P_{\mathrm{H_{2}}}$. If the long-range interaction potential between $\mathrm{Sr}$ and $\mathrm{H_{2}}$ is given by $C_{\mathrm{Sr-H_{2}}}/r^{6}$ (where $r$ is the distance between $\mathrm{Sr}$ and $\mathrm{H_{2}}$), $\gamma$ can be expressed as [36]
@@
Here, by using $C_{\mathrm{Sr-H_{2}}}=166(17)\,E_{h}a_{0}^{6}$ from Ref. [38], where $E_{h}$ is the Hartree energy and $a_{0}$ is the Bohr radius, we obtain
@@ Appendix B Atomic beam flux
We consider determining the atomic beam flux from the oven by measuring the absorption profile of the atomic beam emitted from the oven as shown in Fig. 4. The arguments presented in this section are primarily based on Ref. [35].
@@
$\mathrm{OD}(\nu)=-\ln\frac{I(\nu)}{I_{\text{0}}},$ (5)
where $\nu$ is the laser frequency, $I_{\text{0}}$ is the incident laser intensity, and $I(\nu)$ is the transmitted intensity measured by a photodetector. The OD as a function of frequency $\nu$ can be expressed as
@@
$\mathrm{OD}_{i}(\nu)$ $=nr_{i}\sigma_{0}l_{\text{int}}\frac{1}{\sqrt{2\pi\sigma_{t}^{2}}}\int_{-\infty}^{\infty}\frac{\exp\left(-\frac{v_{t}^{2}}{2\sigma_{t}^{2}}\right)}{1+4\left(\frac{\nu-\nu_{0}^{i}+v_{t}/\lambda}{\Gamma/2\pi}\right)^{2}}dv_{t},$ (7)
where $n$ is the atomic beam density, $r_{i}$ is the isotopic abundance of isotope $i$ for bosons ($^{84}\mathrm{Sr},{}^{86}\mathrm{Sr},{}^{88}\mathrm{Sr}$) and the relative contribution of each hyperfine line $i$ for fermion ($^{87}\mathrm{Sr}$), $\sigma_{0}$ is the resonant scattering cross section without Doppler broadening $\left[\sigma_{0}=3\lambda^{2}/(2\pi)\right]$, $l_{\text{int}}$ is the thickness of the atomic beam along the probe direction (set to $l_{\text{int}}=5\,\mathrm{mm}$ from the aperture size), $\sigma_{t}$ is the transverse Doppler width, $v_{t}$ is the transverse atomic velocity, $\nu_{0}^{i}$ is the resonance frequency of isotope $i$ for bosons and of each hyperfine line $i$ for fermion, $\lambda$ is the probe wavelength ($461\,\mathrm{nm}$), and $\Gamma$ is the FWHM of the resonance transition ($2\pi\times 30.2\,\mathrm{MHz}$). Details on isotope shifts and relative transition strength of $^{87}\mathrm{Sr}$ can be found in Refs. [39, 40]. The absorption profile of the atomic beam is then fitted using Eqs. (6) and (7), with $n$ and $\sigma_{t}$ as fitting parameters (see Fig. 4). From the fitting results, the atomic beam density $n$ is obtained.
The atomic beam flux $\mathcal{F}$ ($\mathrm{atoms\,s^{-1}\,cm^{-2}}$) can be written as
$\mathcal{F}=n\bar{v}_{\text{beam}},$ (8)
where $\bar{v}_{\text{beam}}$ is the mean thermal velocity of the atomic beam, expressed as
$\bar{v}_{\text{beam}}=\sqrt{\frac{8k_{B}T}{\pi m_{\mathrm{Sr}}}}.$ (9)
Using the atomic beam cross-sectional area $A=\pi(l_{\text{int}}/2)^{2}$, the atomic beam intensity, $J$ ($\mathrm{atoms\,s^{-1}\,sr^{-1}}$), is given by
@@
where $\Omega_{\text{aperture}}$ is the solid angle subtended by the aperture as seen from the oven, expressed as
@@
with $L_{\text{aperture}}$ denoting the distance from the oven to the aperture ($140\,\mathrm{mm}$).
On the other hand, the theoretical expression for the atomic beam intensity $J$ in the limit of no collisions occurring in the capillary tubes (Knudsen regime) is written as [35]
$J=\frac{\left(\pi a^{2}\right)\bar{v}_{\text{beam}}n_{\text{oven}}}{4\pi}N_{\text{cap}},$ (12)
where $a$ is the capillary inner radius ($2a=0.3\,\mathrm{mm}$), $N_{\text{cap}}=130$ is the number of capillaries, and $n_{\text{oven}}$ is the atomic density of Sr in the oven. The atomic density is given by $n_{\text{oven}}=P/(k_{B}T_{\text{oven}})$, where $T_{\text{oven}}$ is the atomic temperature inside the oven, and $P$ is the Sr vapor pressure. The latter can be expressed as [41]
$\log_{10}\left(P(\mathrm{Pa})\right)=14.232-\frac{8572}{T_{\text{oven}}\text{(K)}}-1.1926\log_{10}\left(T_{\text{oven}}\text{(K)}\right).$ (13)
@@ Appendix D Dependence of the trapped atom number on the loading rate
where $\gamma$ is the loss coefficient due to background gas collisions, and $\tilde{\beta}$ is the loss coefficient due to two-body collisions. Here, we consider the steady-state trapped atom number $N_{0}$ in two limiting cases: (i) when background gas collisions dominate, and (ii) when two-body collisions dominate.
@@
Thus, $N_{0}$ is proportional to the square root of the loading rate of atoms into the MOT.
According to Ref. [22], in the constant volume regime, the two-body loss coefficient $\beta$ ($\mathrm{cm^{3}\,s^{-1}}$), normalized to the density, can be expressed as
$\beta=\tilde{\beta}\left(\sqrt{2\pi}a\right)^{3},$ (21)
where $a$ is the $1/e$ radius of the density distribution (MOT density $n=n_{0}\exp\left[-\left(r/a\right)^{2}\right]$, with $r$ being the distance from the MOT center and $n_{0}$ the central density).
[…]
We thank Dr. H. Hachisu at the National Institute of Information and Communications Technology, Japan, for providing us with stainless-steel capillary tubes, and M. Mori for his contributions to the experiments.
[…]
This work was supported by JSPS KAKENHI Grant Numbers 23K20849 and 22KJ1163.

### v2

@@ Appendix A Dependence of the loss rate of trapped atoms on background gas pressure
where $R$ is the loading rate of trapped atoms and $\gamma$ denotes the loss coefficient due to collisions with background gas. The dominant component of the background gas is hydrogen ($\mathrm{H_{2}}$). In the following, we assume that the background gas consists solely of $\mathrm{H_{2}}$. Here, $\gamma$ is proportional to the partial pressure of hydrogen $P_{\mathrm{H_{2}}}$. If the long-range interaction potential between $\mathrm{Sr}$ and $\mathrm{H_{2}}$ is given by $C_{\mathrm{Sr-H_{2}}}/r^{6}$ (where $r$ is the distance between $\mathrm{Sr}$ and $\mathrm{H_{2}}$), $\gamma$ can be expressed as [T.Arpornthip2012]
@@
Here, by using $C_{\mathrm{Sr-H_{2}}}=166(17)\,E_{h}a_{0}^{6}$ from Ref. [M.Abdel-Hafiz2019], where $E_{h}$ is the Hartree energy and $a_{0}$ is the Bohr radius, we obtain
@@ Appendix B Atomic beam flux
We determine the atomic beam flux from the oven by measuring the absorption profile of the atomic beam emitted from the oven as shown in Fig. 4. The arguments presented in this section are primarily based on Ref. [M.Schioppo2012].
@@
$\mathrm{OD}(\nu)=-\ln\!\left(\frac{I(\nu)}{I_{0}}\right),$ (5)
where $\nu$ is the laser frequency, $I_{\text{0}}$ is the incident laser intensity, and $I(\nu)$ is the transmitted intensity measured by a photodetector. The OD as a function of frequency $\nu$ can be expressed as a sum of Voigt profiles corresponding to individual isotopic and hyperfine components,
@@
$\mathrm{OD}_{i}(\nu)=\frac{nr_{i}\sigma_{0}l_{\text{int}}}{\sqrt{2\pi}\,\sigma_{t}}\int_{-\infty}^{\infty}\frac{\exp\!\left(-\frac{v_{t}^{2}}{2\sigma_{t}^{2}}\right)}{1+4\left(\frac{\nu-\nu_{0}^{i}+v_{t}/\lambda}{\Gamma/2\pi}\right)^{2}}\,dv_{t},$ (7)
where $n$ is the atomic beam density, $r_{i}$ is the isotopic abundance of isotope $i$ for bosons ($^{84}\mathrm{Sr},{}^{86}\mathrm{Sr},{}^{88}\mathrm{Sr}$) and the relative contribution of each hyperfine line $i$ for fermion ($^{87}\mathrm{Sr}$), $\sigma_{0}$ is the resonant scattering cross section without Doppler broadening $\left[\sigma_{0}=3\lambda^{2}/(2\pi)\right]$, $l_{\text{int}}$ is the thickness of the atomic beam along the probe direction (set to $l_{\text{int}}=5\,\mathrm{mm}$ from the aperture size), $\sigma_{t}$ is the transverse Doppler width, $v_{t}$ is the transverse atomic velocity, $\nu_{0}^{i}$ is the resonance frequency of isotope $i$ for bosons and of each hyperfine line $i$ for fermion, $\lambda$ is the probe wavelength ($461\,\mathrm{nm}$), and $\Gamma$ is the natural linewidth (FWHM) of the transition ($2\pi\times 30.2\,\mathrm{MHz}$). Details on isotope shifts and relative transition strength of $^{87}\mathrm{Sr}$ can be found in Refs. [NIST, S.Mauger2007]. The absorption profile of the atomic beam is then fitted using Eqs. (6) and (7), with $n$ and $\sigma_{t}$ as fitting parameters (see Fig. 4). From the fitting results, the atomic beam density $n$ is obtained.
The atomic beam flux $\mathcal{F}$ ($\mathrm{atoms\,s^{-1}\,cm^{-2}}$) is given by
$\mathcal{F}=nv_{\text{beam}},$ (8)
where $v_{\text{beam}}$ is the mean speed of an effusive atomic beam [H.Metcalf1999]
$v_{\text{beam}}=\sqrt{\frac{9\pi k_{B}T_{\text{oven}}}{8m_{\mathrm{Sr}}}},$ (9)
with $T_{\text{oven}}$ denoting the temperature of the atomic vapor in the oven. Using the atomic beam cross-sectional area $A=\pi(l_{\text{int}}/2)^{2}$, the atomic beam intensity $J$ ($\mathrm{atoms\,s^{-1}\,sr^{-1}}$) is written as
@@
where $\Omega_{\text{aperture}}$ is the solid angle subtended by the aperture as seen from the oven. For small divergence angles it can be approximated by
@@
with $L_{\text{aperture}}$ the distance from the oven to the aperture ($140\,\mathrm{mm}$).
On the other hand, the theoretical expression for the atomic beam intensity $J$ in the limit of no collisions occurring in the capillary tubes (Knudsen regime) is written as [M.Schioppo2012]
$J=\frac{\left(\pi a^{2}\right)\bar{v}n_{\text{oven}}}{4\pi}N_{\text{cap}},$ (12)
where $\bar{v}=\sqrt{8k_{B}T_{\text{oven}}/\pi m_{\text{Sr}}}$ is the mean atomic speed in the oven, $a$ is the capillary inner radius ($2a=0.3\,\mathrm{mm}$), $N_{\text{cap}}=130$ is the number of capillaries, and $n_{\text{oven}}$ is the atomic density of Sr in the oven. The atomic density is given by $n_{\text{oven}}=P/(k_{B}T_{\text{oven}})$, where $P$ is the Sr vapor pressure [L.R.Lide2003]:
$\log_{10}\left(P/\mathrm{Pa}\right)=14.232-\frac{8572}{(T_{\text{oven}}/\mathrm{K})}$
$-1.1926\log_{10}\left(T_{\text{oven}}/\mathrm{K}\right).$ (13)
@@
In the collision-free regime the total flow rate of atoms emitted by the oven at $T_{\text{oven}}=400\,\mathrm{{}^{\circ}C}$ is
$\frac{dN_{\text{oven}}}{dt}=\frac{(\pi a^{2})\bar{v}n_{\text{oven}}N_{\text{cap}}W}{4}\simeq 5.3\times 10^{13}\mathrm{atoms/s},$ (14)
where $W=(8a/3L)/(1+8a/3L)$ the capillary length is the transmission probability (Clausing factor) of the capillaries, with $L=10\,\mathrm{mm}$ the capillary length.
@@ Appendix D Dependence of the trapped atom number on the loading rate
where $\gamma$ is the loss rate due to background gas collisions, and $\tilde{\beta}$ is the effective two-body loss coefficient ($\tilde{\beta}N$ is the effective per-atom loss rate due to two-body collisions). Here, we consider the steady-state trapped atom number $N_{0}$ in two limiting cases: (i) when background gas collisions dominate, and (ii) when two-body collisions dominate.
@@
Thus, $N_{0}$ is proportional to the square root of the loading rate of atoms into the MOT. The fitting result of Fig. 9 yields $\tilde{\beta}=1.11(3)\times 10^{-7}\,\mathrm{s^{-1}}$.
According to Ref. [T.Dinneen1999], in constant-volume regime, the density-normalized two-body loss coefficient $\beta$ ($\mathrm{cm^{3}\,s^{-1}}$) is related to $\tilde{\beta}$ as
$\beta=\tilde{\beta}\left(\sqrt{2\pi}r_{e}\right)^{3},$ (22)
where $r_{e}$ is the $1/e$ radius of the MOT density distribution, given by $n=n_{0}\exp\!\left[-\left(r/r_{e}\right)^{2}\right]$, with $r$ being the distance from the MOT center and $n_{0}$ the central density.